OpenStudy (jimbo69):

Solve. -2/5n= -30 n =

TheSmartOne (thesmartone):

Hi, welcome to OpenStudy! :D

TheSmartOne (thesmartone):

Multiply both sides by -5/2 so that we can isolate n. :)

OpenStudy (jimbo69):

hello! do you think you can help me?

OpenStudy (anonymous):

http://www.cymath.com/answer.php?q=-2%2F5n%3D%20-30%20 this should help you understand better :DDDD

TheSmartOne (thesmartone):

Sure! @Jimbo69

TheSmartOne (thesmartone):

So we need to multiply both sides of the equation by a number so that way we can isolate n, and get our final answer. :)

TheSmartOne (thesmartone):

And so @Jimbo69 like I said earlier, if we multiply both sides by -5/2 we can isolate n. So what is \(\sf\Large \frac{-2}{5} \times \frac{-5}{2} = ?\) AND \(\sf\Large -30\times \frac{-5}{2} = ?\)

OpenStudy (jimbo69):

hold on i'm writing things down

TheSmartOne (thesmartone):

ok :)

OpenStudy (jimbo69):

the first one i don't get but i think the second one is 5/2

TheSmartOne (thesmartone):

nope, that's wrong :/

TheSmartOne (thesmartone):

\(\sf\Large \frac{-2}{5} \times \frac{-5}{2} = ?\) What is 2 x 5 = ?

OpenStudy (jimbo69):

10

TheSmartOne (thesmartone):

and 5 x 2 = 10 so what is 10/10 = ?

OpenStudy (jimbo69):

1

TheSmartOne (thesmartone):

good, now for the next one :)

TheSmartOne (thesmartone):

\(\sf\Large -30\times \frac{-5}{2} = ?\) What is 30 x 5 = ?

OpenStudy (jimbo69):

150

TheSmartOne (thesmartone):

and what is 150/2 = ??

OpenStudy (jimbo69):

75

TheSmartOne (thesmartone):

and that's your final answer :D

OpenStudy (jimbo69):

ok thank you!

TheSmartOne (thesmartone):

Anytime! :D